# Good News for Polynomial Root-finding

### v35

@@ 15.2 Bit operation cost of our e/i tests and root-finders
Apply Thm. 15.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 11 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
Based on [65, Alg. 5.3] and Eqn. (15.2) we obtain a little larger but still reasonably small upper bound
@@
In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (15.3) and (15.4) by a factor of $b$. As by-product we ensure $\epsilon$-isolation and arrive at our Main Theorems 2, 3, and 4.
@@
16 Overview
Classical subdivision iterations converge to all zeros of a polynomial $p$ with linear convergences rate (see Sec. 1.6), but clearly this is very slow when they process a component $\mathbb{C}$ made up of a single suspect square or up to four of them sharing a vertex. In such cases the algorithms of [124, 97] dramatically accelerate convergence to superlinear rate by means of Newton’s or Schröder’s iterations (see Fig. 7), but the authors of [124, 97] have not controlled the bit operation cost of their algorithms.
@@ 16 Overview
Alternatively, one can apply algorithms of [130] to split $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside a disc covering the isolated component $\mathbb{C}$, respectively, and then apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$. Splitting of $p(x)$ into factors, however, does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$. (E.g., splitting $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.) [^42] The near optimal root-finder of [92, 99] relies on by splitting $p(x)$ into factors of comparable degrees, but under the assumption of computing with a rather high precision of order $d\log(d)$. These problems disappear if one of the factors, say, $g$ has a small degree: one can deflate it, approximate its zeros, and then approximate the zeros of $f(x)=p(x)/g(x)$ by applying implicit deflation (cf. (1.14)).
We stay within subdivision framework and like [124, 97, 22, 23] partition the union of all suspect squares at each subdivision step into pairwise isolated connected components but propose a novel algorithm that compresses such a component $\mathbb{C}$ into a near minimal disc containing all roots that lie in the component.
This algorithm enables us to decrease the estimates of Secs. 8.3 and 15.2 for the overall complexity of root-finding by this factor, thus providing formal support for our Main Theorems 2-4.
We organize our presentation in this part as follows. In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components. In Sec. 18 we devise and analyze our first compression algorithm. In particular we observe in Sec. 18.4 that the whole subdivision process involves at most $12m-$ 1 compression steps. In Sec. 19 we prove that in addition it only involves $O(m)$ e/i steps. Together with our estimates for the complexity of an e/i step this enables us to bound the overall complexity of our root-finders and reach near optimal bound on the bit operation complexity of our root-finders. We reach that bound in Sec. 20 by devising and analyzing our alternative compression algorithm. In that section we also present an alternative fast root-counter in an isolated disc.
We use some results of Part I, Alg. 5.1 of Part II, and Sec. 10 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained. Our fast root-counter of Sec. 20.4 can serve as alternative to root-counters of Part II.
@@ 17 Component tree and compact components
17.1 Component tree: definition
Partition the union of all suspect squares at a subdivision step into components; a step can break some of them into more than one component.
Represent all such partitions with a component tree $\mathbb{T}$.
At most $m$ components, said to be the leaves of the tree, can be covered by isolated discs of radius at most $\epsilon$ for a fixed tolerance $\epsilon\geq 0$. The other vertices are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree $\mathbb{T}$. The edges of the tree connect parents and children. The tree does not shows modifications of components between their partitions represented by the edges.
Observation 17.1.
Having at most $m$ leaves, the tree $\mathbb{T}$ has at most $m-1$ other nodes. These upper bounds are sharp – attained where a subdivision process defines the complete binary tree $\mathbb{T}$ with $m$ leaves.
17.2 Edge length, non-adjacent suspect squares, and compact components
The edges of the tree $\mathbb{T}$ represent subdivision steps between partitions of components. Cor. 17.1 will bound the number of such steps; one can view this as the edge length.
@@ 17.2 Edge length, non-adjacent suspect squares, and compact components
Observation 17.2.
@@ 18.1 Compression algorithm: high level description
Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.4) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ to scaling by $\rho$. Observation 17.3 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
@@
INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^43] Newton’s and Schröder’s iterations of [124, 97] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while (18.1) only requires 3.5-isolation.
@@
OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$.
@@
COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, $\rho^{\prime}$ such that
@@ 18.2 Correctness of compression algorithm
Now (18.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 18.1.
Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a zero $z_{g}$ of $p(x)$. Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
@@
Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 18.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
@@
Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 18.1.
Next wlog assume that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
@@
This implies claim (ii) for the output of Alg. 18.1 since $\rho^{\prime}>\epsilon$.
@@ 18.3 Elaboration upon the compression algorithm
We can readily compute $m=\#(D(0,1))$ for isolated input disc $D(0,1)$ of Alg. 18.1, e.g., by applying Alg. 3.1, and can extend this recipe to any disc $D(c,\rho)$ by applying map (1.4) and Prop. 3.1. Moreover, in Sec. 20.4 we describe a very fast alternative root-counter.
Furthermore, based on Prop. 3.2 we can compute close approximations $\tilde{s}_{1}$ to the sum $s_{1}$ of the $m$ zeros of $p$ lying in that disc and then closely approximate their average $c^{\prime}:=s_{1}/m$, lying in the convex hull $CH(D(c,\rho))$. Finally, we can approximate the $m$th smallest root radius $r_{d-m+1}(c^{\prime},p)$ by applying our root radius estimation of Sec. 5, based on $\ell$-tests for $\ell=m$. There we can apply $\ell$-tests of Part II or combine e/i tests of Sec. 11 with reduction of $m$-test to e/i tests in Sec. 12.
Computation of a sufficiently close approximation to the sum $s_{1}$ of the zeros of $p$ lying in the disc $D(0,1)$ and of the average $s_{1}/m$ is too costly to support Thms. 1.1 and 1.3 (we would need to evaluate $p(x)$ at $O(b\log(d)$ points and similarly for bit operation cost, but we fix this deficiency in Sec. 20 by presenting another variant of compression algorithm.
@@ 18.4 Counting $\epsilon$-compression steps
$c^{\prime}$ approximates within $\epsilon$ all $m$ roots lying in the disc $D(0,1)$ if $\rho^{\prime}\leq\epsilon$ for the output value $\rho^{\prime}$ of Alg. 18.1. Then call the compression final and stop computations.
@@
Extend Lemma 17.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg. 18.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
Until that moment apply no new compression steps and hence apply less than $2m$ such steps overall – no more than there are nodes in the partition tree $\mathbb{T}$.
Now elaborate upon Observation 17.1 as follows.
@@
For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots. Then at most $m$ final compression steps among at most $12m-$ 1 compression steps overall output isolated discs of radii at most $\epsilon$ that cover all roots lying in $C$.
Corollary 18.1.
Under the assumptions of Thm. 18.1 all subdivision steps process at most $36m-$ 3 compact components, made up of at most $1224m-$ 12 suspect squares; this involves at most $1224m-$ 12 e/i tests.
@@
A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a compression step or follow a non-final compression step, that is, such components are processed in at most $36m-$ 3 subdivision , which involve at most $1224m-$ 12 e/i tests overall. ∎
@@ 19.1 An upper estimate
The upper estimate $12m-$ 1 of Thm. 18.1 is sharp – reached in the case of the complete binary tree $\mathbb{T}$ and implying that our accelerated subdivision process can involve $12m-$ 1 suspect squares (e/i steps). This reaches lower estimate up to a constant factor:
@@ 19.2 The first lemma
Our proof of Thm. 19.1 begins with
@@ 19.3 The second lemma
Fix two integers $i\geq 0$ and $g\geq 3$, a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.2. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
@@
By virtue of Cor. 18.1 all compact components together involve $O(m)$ suspect squares and e/i tests; it remains to prove that we only involve $O(m)$ other e/i tests, for which $n_{j}\geq 2$ for all $j$.
Next we extend Lemma 19.2 by allowing compression steps. We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ may perform compression, after which we apply subdivision process to a single suspect square with side length at most $4\Delta_{j}$.
The next two subdivision steps perform only five e/i tests, that is, at most $510m-$ 5 e/i tests are performed at all pairs of subdivision steps following all (at most $12m-$ 1) compression steps.
Then again, we can exclude all these e/i tests and bound by $O(m)$ the number of e/i tests in the remaining subdivision process, which is not interrupted with compression steps anymore.
Furthermore, we resume subdivision process after a compression step with a suspect square whose side length decreased (in two discounted subdivision steps) from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
@@ 19.4 Counting e/i tests at partition steps
Let $h$ be the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.

### v36

@@ 18.2 Bit operation cost of our e/i tests and root-finders
Apply Thm. 18.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 14 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
Based on [68, Alg. 5.3] and Eqn. (18.2) we obtain a little larger but still reasonably small upper bound
@@
In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (18.3) and (18.4) by a factor of $b$, and arrive at our Main Theorems 2, 3, and 4.
@@
19 Overview and two alternative approaches
Classical subdivision iterations converge to the roots with linear convergence rate (see Sec. 2), which is slow when many consecutive iterations process a component $C$ made up of a single suspect square or up to four of them sharing a vertex. In this case the algorithms of [128, 101] dramatically accelerate convergence by means of applying Newton’s or Schröder’s iterations (see Fig. 8), which converge with superlinear rate to the convex hull of all roots lying in $C$. The diameter of the convex hull, however, can be too large, and then [128, 101] apply subdivision iterations again.
The resulting root-finder recursively intertwines subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder. [128, 101] have only estimated the number of ops but not of the bit operations involved. [23, 24] estimated the bit operation complexity, but only under a very strong assumption about the output (see Remark 4.2 and Sec. 12.4).
@@ 19 Overview and two alternative approaches
Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [134], which splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively. Then one can apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$. Splitting of $p(x)$ into factors does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$. (E.g., split $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.)
Overall asymptotic bit-complexity for high precision root-finding can still be near-optimal if a root-finder recursively splits $p(x)$ into factors of comparable degrees. This is achieved in [96, 103], but the resulting root-finder is much involved and extensively uses the coefficients of $p$.
The problems caused by splitting disappear if one of the factors, say, $g$ has a small degree; then we can deflate it, approximate its zeros readily, and approximate the zeros of $f(x)=p(x)/g(x)$ by applying to it a black box root-finder and using implicit deflation based on (3.3) or the equation NIR$(p)=$NIR(f)+limit-$(f)+$NIR$(g)$ for $p=fg$.
One would lose optimality, however, by always deflating low-degree factors: to deflate $k$ times one would need at least order of $bkmd$ bit operations overall, that is, would exceed the lower bound $0.25(d+1)db$ of Cor. 24.2 by at least a factor of $k$.
We stay with subdivision framework but instead of Newton’s or Schröder’s iterations, used in [128, 101], apply our novel algorithm that compresses a component isolated from its external roots into a near-minimal disc that keeps all internal roots. This enables us to decrease the estimates of Secs. 12.3 and 18.2 for the overall complexity of root-finding by a factor of $b$, thus completing proof of our Main Theorems 2-4.
We organize our presentation in this part as follows. In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components. In Sec. 21 we devise and analyze our first compression algorithm. In particular we observe in Sec. 21.4 that the whole subdivision process involves at most $12m-$ 1 compression steps, reduced to $O(m)$ e/i tests. In Sec. 22 we prove that in addition it also involves $O(m)$ e/i steps. We reach a near-optimal bound on the bit operation complexity of our root-finders in Sec. 23 by devising and analyzing our compression algorithm and recalling our estimates for the complexity of an e/i step. In that section we also present a fast root-counter in an isolated disc.
We use some results of Part I, Alg. 9.1 of Part II, and Sec. 13 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained.
@@ 20 Component tree and compact components
20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
Recall the definition of the component tree $\mathbb{T}$ of Sec. 4, which represent partition of the union of suspect squares at every subdivision step and representation of such partition with the component tree $\mathbb{T}$ of Sec. 4. The tree has at most $m$ components-leaves that represent the output $\epsilon$-discs and at most $m-1$ other vertices, which represent components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree $\mathbb{T}$. The edges of the tree connect parents and children, representing subdivision steps between partitions of components. Cor. 20.1 below bounds the number of such steps; one can consider this number the edge length.
@@ 20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
Figure 9: Two discs superscribing two non-adjacent suspect squares are not minimal but do not overlap.
Observation 20.1.
@@ 21.1 Compression algorithm: high level description
Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.3) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ by scaling by $\rho$. Observation 20.2 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
@@
INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^34] Newton’s and Schröder’s iterations of [128, 101] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while Alg. 21.1 only requires 3.5-isolation (see (21.1)).
@@
OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$. [Hence the disc $D(c^{\prime},\rho^{\prime})$ is 5/8-rigid.]
@@
COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, and $\rho^{\prime}$ such that
@@ 21.2 Correctness of compression algorithm
Now (21.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 21.1.
Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a root $z_{g}$. Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
@@
Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 21.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
@@
Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 21.1.
Next assume wlog that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
@@
This implies claim (ii) for the output of Alg. 21.1 since $\rho^{\prime}>\epsilon$.
@@ 21.3 Elaboration upon the compression algorithm
We readily (i) count roots in isolated input disc $D(0,1)$ of Alg. 21.1 by applying the fast root-counter of Sec. 23.4 or alternatively Alg. 7.1 and Prop. 7.1 and (ii) compute a complex point $c^{\prime}$ lying in or near the convex hull of the roots in $C$ by applying fast algorithm of Sec. 23. Finally, we estimate the root radius $\rho^{\prime}:=r_{d-m+1}(c^{\prime})$ by applying Alg. 9.1, which incorporates the $m$-test of of Sec. 23.4, or alternatively Alg. 7.1 and Prop. 7.1.
@@ 21.4 Counting $\epsilon$-compression steps
Recall from Sec. 1.3 that the output disc $D(c^{\prime},\rho^{\prime})$ of Alg. 21.1 is said to be an $\epsilon$-disc if $\rho^{\prime}\leq\epsilon$; in this case call the compression by Alg. 21.1 final and output that disc.
@@
Extend Lemma 20.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg. 21.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
From part (ii) of Sec. 4 recall the bound $22m-$ 2 on the number of compression steps and obtain
@@
For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots. Then the computations involve at most $66m-$ 6 compact components, made up of at most $2424m-$ 24 suspect squares; that is, these computations involve at most $2424m-$ 24 e/i tests.
@@
A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a final compression step or follow a non-final compression step and end with breaking a component. Hence all such components are processed in at most $66m-$ 6 subdivision step, each involving at most 4 e/i tests. ∎
@@ 22.1 An upper estimate
The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$. Each compression step is preceded by an e/i test, and so we perform at least $22m-$ 2 e/i steps. According to the following result, which is the main theorem of this section, this lower bound on the number of e/i tests involved is sharp up to a constant factor:
@@ 22.2 The first lemma
We begin our proof of Thm. 22.1 with
@@ 22.3 The second lemma
Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 20.1. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
@@
Next we extend Lemma 22.2 by allowing compression steps. We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
It remains to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
Furthermore, after a compression step we resume subdivision process with a suspect square whose side length bound decreases in two next subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
@@ 22.4 Counting e/i tests at partition steps
Let $h$ denote the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.